# ViewMask-1-to-3: Multi-View Consistent Image Generation via Multimodal Diffusion Models

### v1

@@ 4.2 Text to Multi-view Images
In the T2MV task, we demonstrate the ability to generate multi-view images of an object directly from text inputs. Models like MV-Adapter [huang2024mvadapter] and TRELLIS [xiang2024structured] support T2MV and I2MV using different checkpoints or architectures. Zero123++ [shi2023zero123++] and ViVid-1-to-3 [kwak2023vivid] generate reference images with a T2I model before extracting the subject for novel views. In contrast, our method directly produces novel views that preserve both the specified viewpoint and the prompt’s semantic content. As shown in Fig. 5, we present examples from the Objaverse dataset that were selected as test samples.
@@ 4.4 Task Extension
Unifed Multimodal Understanding and Generation
In addition to the aforementioned image-to-multiview images(I2MV) and text-to-multi-view images(T2MV) capabilities, the Stage-1 pretraining endows our model with fundamental multimodal understanding (MMU) and text-to-image (T2I) generation abilities. As shown in Fig. 7, the model can perform basic understanding and generation tasks without any additional fine-tuning. Building upon these foundational abilities, our model can naturally extend toward a unified multimodal understanding and generation framework, enabling effortless adaptation to a wide range of new multimodal tasks.
@@
Inspired by Adobe’s Project Turn Style, which enables re-angling and rotating 2D illustrations as if they were 3D objects, we extend this capability in a training-free manner based on our ViewMask-1-to-3 framework. Benefiting from the inherent properties of discrete diffusion, our model can rotate a specified object by a given angle and perform image completion while preserving the original background, all without additional fine-tuning, as these abilities are implicitly embedded during diffusion training. These cases demonstrate that our framework not only extends toward a unified diffusion architecture, but also exhibits strong flexibility and generalization across diverse generation and reasoning tasks.
@@ 5 Conclusion
In this work, we introduce ViewMask-1-to-3, a discrete diffusion framework for multi-view image generation. Experiments on GSO and 3D-FUTURE demonstrate strong performance, highlighting discrete diffusion as an effective and flexible alternative to traditional continuous diffusion and 3D-aware approaches. Our results also suggest the model’s potential for general multimodal understanding (MMU), text-to-image (T2I) generation, text-to-multi-view (T2MV) synthesis, and training-free 2D turn-style image manipulation, pointing to a possible path toward more unified and scalable multi-view generation.
Despite its effectiveness, limitations remain, such as restricted elevation angles, a small number of viewpoints, and differences in lighting during turn-style generation. Future work will explore increasing viewpoint diversity, improving resolution, and better handling lighting variations to further advance unified and scalable multi-view generation.
[…]
Supplementary Material
@@ A.1 Baseline Setting
Figure 8: Perspective vs. Orthographic Projection.
As shown in Fig. 8, most of the baselines we select—such as Zero-1-to-3 [liu2023zero], Zero-1-to-3 XL [liu2023zero, objaverseXL], ViVid-1-to-3 [kwak2023vivid], Zero123++ [shi2023zero123++], EpiDiff [huang2024epidiffenhancingmultiviewsynthesis], AR-1-to-3 [zhang2025ar], TRELLIS [xiang2024structured], and our own method ViewMask-1-to-3 are trained on data rendered with perspective projection. Among all these methods, only MV-Adapter [huang2024mvadapter] is trained using data rendered with orthographic projection.
A.2 Ground-Truth Rendering
@@
Appendix B Additional Qualitative Comparison
@@ B.1 Image to Multi-view Images
Figure 9: Qualitative results on T2MV. Examples of multi-view object generation from text prompts using MV-Adapter, Trellis, ViewMask-1-to-3.
@@
B.2 Text to Multi-view Images
To further illustrate the qualitative performance differences, we visualize and compare the results produced by MV-Adapter [huang2024mvadapter], TRELLIS [xiang2024structured], and our method, as shown in Fig. 9. In contrast to MV-Adapter, whose orthographic projection setup limits the amount of recoverable geometric detail, our method produces richer visual details across views. Compared with TRELLIS, which represents one of the strongest baselines, our approach achieves comparable visual quality while requiring only a single inference pass and using the same unified checkpoint across all tasks.
B.3 Task Extension
@@
During inference, ViewMask employs task-specific input templates that leverage discrete tokens for both vision and language modalities. For multimodal understanding (MMU) tasks, the input sequence is structudarkred as:
@@
Training-Free 2D Turn-Style
As shown in Fig. 7, ViewMask-1-to-3 implements a similar training-free capability, allowing 2D illustrations to be re-angled as if they were 3D. Our pipeline performs two sequential inference passes to achieve multi-view image generation while preserving the original scene context. In the first pass, the model receives the object obtained from the original image and generates its appearance across multiple viewpoints. In the second pass, the object is masked and provided to the model again to reconstruct the masked regions, restoring the original background while integrating the previously generated object views. By fusing the outputs of both stages, the final results yield consistent multi-view objects seamlessly incorporated into the original scene, all achieved without additional fine-tuning.

### v2

@@ 4.2 Text to Multi-view Images
In the T2MV task, we demonstrate the ability to generate multi-view images of an object directly from text inputs. Models like MV-Adapter [18] and TRELLIS [46] support T2MV and I2MV using different checkpoints or architectures. Zero123++ [39] and ViVid-1-to-3 [22] generate reference images with a T2I model before extracting the subject for novel views. In contrast, our method directly produces novel views that preserve both the specified viewpoint and the prompt’s semantic content. As shown in Fig. 5, we present representative test samples from the Objaverse dataset and compare qualitative results with MV-Adapter [18], TRELLIS [46], and SPAD [20], showing that our method achieves comparable visual quality and consistency.
@@ 4.4 Task Extension
Unified Multimodal Understanding and Generation
In addition to the aforementioned I2MV and T2MV capabilities, the Stage-1 pretraining endows our model with fundamental multimodal understanding and text-to-image generation abilities. As shown in Fig. 7, the model can perform basic understanding and generation tasks without any additional fine-tuning. Building on these abilities, our model naturally extends toward unified multimodal understanding and generation, evaluated on GenEval [13] for compositional text-to-image generation. ViewMask-1-to-3 achieves the best overall score (0.72), outperforming unified models Show-o (0.68) [49] and MMaDA (0.63) [52], and surpassing generation-only models like DALL·E 3 (0.67) [3], comparable to SANA-1.5 (0.72) [48], as shown in Table 4. Compared with the architecturally similar MMaDA, ViewMask-1-to-3 achieves a 14% overall improvement (0.72 vs. 0.63). This significant gain highlights the practical potential of discrete diffusion architectures.
@@
Inspired by Adobe’s Project Turn Style, which enables re-angling and rotating 2D illustrations as if they were 3D objects, we extend this capability in a training-free manner based on our ViewMask-1-to-3 framework. Benefiting from the inherent properties of discrete diffusion, our model can rotate a specified object by a given angle and perform image completion while preserving the original background, all without additional fine-tuning, as these abilities are implicitly embedded during diffusion training. These cases show that our framework extends to a unified diffusion architecture while maintaining strong flexibility and generalization across diverse tasks.
@@ 5 Conclusion
In this work, we introduce ViewMask-1-to-3, a discrete diffusion framework for multi-view image generation. Experiments on GSO and 3D-FUTURE demonstrate its strong performance, showcasing discrete diffusion as an effective and flexible approach for multi-view synthesis. Our results also highlight its potential for MMU, T2I, T2MV, and training-free 2D turn-style, paving the way toward more unified and scalable multi-view generation.
Despite strong performance, there is room to expand viewpoint diversity and improve lighting robustness. Future work will focus on enhancing scalability, resolution, and generalization for more unified multi-view generation.
Acknowledgements
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000.
@@ A.1 Baseline Setting
Most of the baselines we select—such as Zero-1-to-3 [28], Zero-1-to-3 XL [28, 7], ViVid-1-to-3 [22], Zero123++ [39], EpiDiff [19], AR-1-to-3 [57], TRELLIS [46], and our own method ViewMask-1-to-3 are trained on data rendered with perspective projection. Among all these methods, only MV-Adapter [18] is trained using data rendered with orthographic projection.
A.2 Experimental Settings
[…]
A.3 Ground-Truth Rendering
@@
Table 4: Evaluation on Image Generation Benchmarks (GenEval). Our model achieves strong performance in unified models and is also comparable to methods among generation-only models.
Model Architecture Single Obj. Two Obj. Counting Colors Position Color Attr. Overall
Generation-Only
SDv1.5 [38] Diffusion 0.97 0.38 0.35 0.76 0.04 0.06 0.43
SDv2.1 [38] Diffusion 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E 2 [37] - 0.94 0.66 0.49 0.77 0.10 0.19 0.52
DALL-E 3 [3] - 0.96 0.87 0.47 0.83 0.43 0.45 0.67
LlamaGen [41] AR 0.71 0.34 0.21 0.58 0.07 0.04 0.32
FLUX.1 [Dev] [23] Diffusion 0.98 0.81 0.74 0.79 0.22 0.45 0.66
SDXL [36] Diffusion 0.98 0.74 0.39 0.85 0.15 0.23 0.55
OmniGen [47] Diffusion 0.98 0.84 0.66 0.74 0.40 0.43 0.68
SANA-1.5 [48] Diffusion 0.99 0.85 0.77 0.87 0.34 0.54 0.72
Unified Understanding & Generation
SEED-X [12] AR 0.97 0.58 0.26 0.80 0.19 0.14 0.49
LWM [26] AR 0.93 0.41 0.46 0.79 0.09 0.15 0.47
Janus [45] AR 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Show-o [49] AR+Diff. 0.98 0.80 0.66 0.84 0.31 0.50 0.68
VAR-GPT [59] AR 0.96 0.53 0.48 0.83 0.13 0.21 0.53
MMaDA [52] Discrete Diff. 0.99 0.76 0.61 0.84 0.20 0.37 0.63
[…]
Appendix B Additional Results
@@ B.1 Image to Multi-view Images
B.2 Task Extension
@@
During inference, ViewMask-1-to-3 employs task-specific input templates that leverage discrete tokens for both vision and language modalities. For multimodal understanding (MMU) tasks, the input sequence is structured as:
@@
We quantify the text-to-image generation capability of our model on GenEval, which evaluates object-centric generation under compositional prompts with diverse object attributes. Table 4 shows that ViewMask-1-to-3 achieves a strong overall score of 0.72, outperforming the architecturally similar MMaDA by 14% (0.72 vs. 0.63), highlighting the practical potential of discrete diffusion architectures.